A missile is launched with a velocity less than the escape velocity the sum of its kinetic and potential energy is

Grade:12th pass

4 Answers

Vikas TU
7 years ago
Kinetic energy would be obvious = 0.5mv^2
and the potential Energy would be= -Gm/r^2
Now total energy = 0.5mv^2 – Gm/r^2 hence would be negative as r approaches infinity.
Siddharth Lathi
5 years ago
Negative. For the velocity than escape velocity, the missle is bound due to the gravtitaion field of the earth. Hence its total energy is negative
Ashish Mishra
5 years ago
Total energy is conserved, hence for the complete motion total energy is k.e+ P.e.
K.e = -p.e/2=-t.e( total energy).
T.e is negative.
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Kinetic energy would be obvious = 0.5mv^2
and the potential Energy would be= -Gm/r^2
Now total energy = 0.5mv^2 – Gm/r^2 hence would be negative as r approaches infinity.

Thanks and Regards